OpenStudy (anonymous):

this the question first estimate the answer.Then multiply to find the exact answer. simplify all the answers. 4 1/3*2 3/4

OpenStudy (anonymous):

can some one help, me with this problem

OpenStudy (espex):

Well the first part of the question asks you to estimate the answer, so if you start with \[4\frac{1}{3}\] and you have \[2\frac{3}{4}\] of them, 2.75 * 4 is 8 + just over half again will give you 8 + 2(and a little) so the whole number is about 10ish. Next you have the 2.75 * 1/3, which is 2/3 + just over half again, 2/3 + 1/3(and a little) is 3/3. Add the two together and you have 10ish+3/3ish = 11ish. Next you need to multiply the numbers out and simplify the answer. If you're doing it by hand, start by making improper fractions: \[4\frac{1}{3}\times 2\frac{3}{4} \rightarrow \frac{13}{3}\times \frac{11}{4}\] then multiple horizontally: \[\frac{13}{3}\times \frac{11}{4} \rightarrow \frac{143}{12}\] and finally you reduce your answer: \[\frac{143}{12} \rightarrow 132\frac{11}{12} \rightarrow 132.92\]